OpenStudy (anonymous):

how can you find the distance between the pair of points? (-2,-5) and (2,3)

OpenStudy (anonymous):

use distance formula ... simple

OpenStudy (anonymous):

\[\in 2D, dist = \sqrt((x1 - x2)^2 + (y1 - y2)^2)\]

OpenStudy (anonymous):

so the answer is 4 (square root) of 5, right? (my keyboard does not have the symbols)

myininaya (myininaya):

neither does ours we use the equation editor below its fun \[\sqrt{(-2-2)^2+(-5-3)^2}=\sqrt{(-4)^2+(-8)^2)}=\sqrt{16+64}=\sqrt{80}=\sqrt{16*5}=\sqrt{16}\sqrt{5}\] \[=4\sqrt{5}\]